# A Principle of Targeted Intervention for Multi-Agent Reinforcement Learning

### v3

@@ 4 MARL for Sequential Decision Making in MAIDs
We now show how the pre-strategy intervention introduced above is implemented in MARL for sequential decision making. We first show that each Markov game can be transformed into one multi-agent influence diagram (MAID). We model MARL using team reward Markov games littman1994markov . The main idea is establishing a connection that Markov games can be transformed into MAIDs, allowing our framework to apply. The formal definition of a Markov game and the detailed process of representing it as an MAID by matching their respective variables are provided in Appendix G. Building on this foundation, the subsequent subsections will extend the MARL interaction paradigms introduced before to Markov games (Section 4.1), discuss the solvability of MARL learning paradigms under these MARL interaction paradigms (Section 4.2), and specify how the pre-strategy intervention in Eq. (3) is integrated into generic MARL learning algorithms (Section 4.3).
4.1 Extending MARL Interaction Paradigms to Markov Games
$D^{t}_{a}$$U^{t}$$D_{h}^{t}$$S^{t}$$U_{a}^{t}$$U_{h}^{t}$$Z^{t}$$S^{t+1}$
@@ 4.1 Extending MARL Interaction Paradigms to Markov Games
Figure 3: MAIDs and their relevance graphs for dynamic simultaneous-move games. For conciseness, we only demonstrate one transition from timesteps $t$ to $t+1$. The variables used in one-shot decision making (see Figure  1) are decorated with timesteps. The semantics of these variables remain unchanged. The pink circles with $s_{t}$ and $s_{t+1}$ indicate states for timesteps $t$ and $t+1$. The green diamonds with $U^{t}$ indicate the total utility variables. The incoming edge from an individual utility variable ($U^{t}_{h}$ and $U^{t}_{a}$) to the total utility variable indicates that the agent’s individual utility contributes to the team reward.
$D^{t}_{a}$$I^{t}$$D^{t}_{h}$$S^{t}$$U^{t}_{a}$$U^{t}_{h}$$Z^{t}$$U^{t}$$S^{t+1}$
@@
We now demonstrate how to extend the three MARL interaction paradigms (see Definition 3.1) to Markov games through the lens of MAIDs. Each interaction paradigm under sequential decision making in dynamic environments with simultaneous or sequential moves can be extended from one-shot decision making introduced in Section 3.2, by considering timesteps and states. With representation of the total utility variable, we can derive the MAID graphical representation of these three interaction paradigms in Markov games, which are illustrated in Figures 3 and 4. Recall that the variable $\boldsymbol{Z^{t}}$ has different semantic meanings across different MARL interaction paradigms, similar to the variable $\boldsymbol{Z}$ in Definition 3.1.
It can be observed that the widespread global intervention paradigm implemented for MARL algorithms (e.g., applying a centralized coordinator wang2020roma ; na2024lagma ; liu2023lazy ) can be justifiably interpreted with a visualization. More importantly, it opens a new door to design novel MARL interaction paradigms to solve a broad class of problems with the help of MAIDs, such as the targeted intervention paradigm proposed in this paper to solve the team reward Markov games.
@@ 4.2 Solvability of MARL Learning Paradigms under MARL Interaction Paradigms
Independent Learning. Independent learning (IL) algorithms (claus1998dynamics, ) are the direct application of single-agent reinforcement learning algorithms (e.g., temporal-difference (TD) learning (sutton2018reinforcement, )[Ch. 6] and the actor-critic algorithms (konda1999actor, )) to MARL. If a relevance graph is cyclic, it is not guaranteed to reach an NE by solely determining each agent’s decision variables with a generalized backward induction algorithm in analogy to IL, as discussed in Section 2.4. This aligns with the self-organization paradigm (Figure 3(a) and 4(a)), reflecting the non-stationarity dilemma (hernandez2019survey, ). Equipping IL algorithms with the global intervention (Figures 3(b) and 4(b)) and the targeted intervention (Figures 3(c) and 4(c)) paradigms can generate acyclic relevance graphs. Such acyclic relevance graphs can guarantee the tampered IL algorithms reaching an NE with theoretical guarantees.
Centralized Training and Decentralized Execution. As discussed in Remark 2.5, the only way to address cyclic relevance graphs is enabling two agents’ decision making to be matched. One solution is transforming a cyclic relevance graph to a component graph, where each maximal strongly connected component (SCC) is regarded as a supernode de2023graph . More specifically, a maximal SCC consists of all the decision variables forming a cyclic relevance graph at each timestep. Koller and Milch koller2003multi showed that solving the acyclic component graph [^3] A component graph is always acyclic (cormen2022introduction, ). by the generalized backward induction algorithm can reach an NE. This is in coincidence associated with the centralized training and decentralized execution (CTDE) employed to reach the maximum cumulative team rewards in Markov games (oliehoek2008optimal, ; littman2001value, ). In summary, CTDE can be seen as an MARL learning paradigm to address the non-stationarity dilemma of vanilla IL algorithms existing in the self-organization paradigm, rather than switching to other MARL interaction paradigms.
@@ 4.3 Pre-Policy Learning in MARL
As we have established how MARL interaction paradigms can be represented in MAIDs and subsequently how it can fit in the team-reward Markov games, the techniques developed in Section 3.3.2 can guide us to realizing the pre-strategy intervention in the team-reward Markov games. Following Definition 3.2, the requirement is to learn a pre-policy $\delta_{pre}:dom(Pa(D))\times\mathcal{Z}\rightarrow\Delta(dom(\sigma_{pre}))$ to generate the pre-strategy $\sigma_{pre}$ for the possible optimal NEs. This can be realized by maximizing the causal effect expressed in Eq. (3). Our proposed algorithm, which can be integrated with generic MARL algorithms, is referred to as Pre-Strategy Intervention (PSI).
Practically, Eq. (3) is implemented by Monte Carlo sampling, the RHS of which can be reformulated as a posterior distribution over a Nash equilibrium $\sigma$, as follows:
$P(\sigma\mid U_{tot}=u^{*},\operatorname{do}(\sigma_{pre}))\propto P_{\mathcal{M}[\sigma]}(U_{tot}=u^{*}\mid\sigma)P_{\sigma}(\sigma\mid\operatorname{do}(\sigma_{pre})).$ (4)
Maximizing the posterior distribution $P(\sigma\ |\ U_{tot}=u^{*},\operatorname{do}(\sigma_{pre}))$ in Eq. (4) reveals a fact that observing the total utility variable $U_{tot}$ (rewards) and a pre-strategy intervention $\operatorname{do}(\sigma_{pre})$ are the key conditions for seeking the optimal joint agent strategy profile denoted by $\sigma$. This more intuitively illustrates the maximization of the causal effect of the pre-strategy intervention in Eq. (3), which provides a clear probabilistic objective that can be optimized by generic MARL algorithms.
@@ 5.1 Experimental Setups
We evaluate our proposed Pre-Strategy Intervention (PSI) in two environments: the Multi-Agent Particle Environment (MPE) lowe2017multi , a simultaneous-move cooperative navigation game, and the Hanabi card game Bard_2020 , which challenges coordination within a multi-agent system under partial observability. Hanabi is known for having multiple, distinct equilibria, making it an ideal testbed for analyzing convergence to a specific Nash equilibrium. In both settings, the intervention is applied to a single targeted agent to assess its effectiveness in achieving the additional desired outcome and improving the task completion. For the experiments presented in this paper, we select a fixed agent for intervention in each environment, which allows for a controlled analysis of the PSI. Full details of these environments, including specific experimental configurations and how the additional desired outcomes are defined and implemented, are provided in Appendix H. In all environments, intrinsic rewards are used for measuring reachability of the additional desired outcomes, while extrinsic rewards are used for measuring the task completion. Results from 5 random seeds are reported as means with $95\%$ confidence intervals.
Implementation of Our Method. Our PSI is implemented by designing a pre-policy module (a GRU or MLP, matching agent’s backbones) to realize the targeted intervention paradigm, which takes a concatenated input of environmental observations and a measure representing an additional desired outcome (formulated as an intrinsic reward). [^4] This implementation maintains solvability of the PSI (see Appendix F.2). It outputs an embedding, which is then fed to Q-value functions for value-based MARL methods or critics for policy-based methods (see Appendix H for details).
Baselines. All baselines are implemented with the same architecture and training setups as our method. We evaluate our method against several base MARL algorithms referred to as Base MARL (self-organization paradigms): for MPE, these include DQN (mnih2015human, ), VDN (sunehag2017value, ) and QMIX (rashid2020monotonic, ); for Hanabi, we use IPPO (schulman2017proximal, ), MAPPO (yu2022surprising, ) and PQN (gallici2024simplifying, ) (with PQN-IQL for IL and PQN-VDN for CTDE variants). An ablation study that removes the pre-policy module while retaining the intrinsic reward maximization examines the effectiveness of the targeted intervention paradigm without the confounding effect of the pre-policy module. This ablated variant is referred to as Intrinsic Reward that still satisfies the targeted interaction paradigm, with removing pre-decision variables $D_{pre}$ in Figures 3 and 4 and connecting the $Z^{t}$ interpreted as the intrinsic rewards to maximize directly to the decision variables $D$. To emphasize the strength of our targeted intervention paradigm, we also compare against the existing approaches implemented in the global intervention paradigm, modifying LIIR du2019liir or LAIES liu2023lazy as representative policy-based or value-based methods. This baseline is referred to as Global Intervention.
@@ 5.2 Results and Analysis
(a) MPE: Main results (extrinsic left, intrinsic right).
(b) MPE: Targeted vs. Global
(c) Hanabi: Main results (extrinsic left, intrinsic right).
(d) Hanabi: Targeted vs. Global
@@
Figure 5: In Subfigures a and c, our method PSI is compared against Intrinsic Reward and ase MARL; these plots show both extrinsic (task completion) and intrinsic (additional desired outcome) returns. Our targeted intervention paradigm (PSI) is compared against the methods of global intervention paradigm in Subfigures b and d with extrinsic returns.
Additional Goals Improving Task Performance. As Figures 5(a) and 5(c) show, both our PSI and the Intrinsic Reward baseline outperform base MARL algorithms. These results confirm that incorporating guidance from an additional desired outcome enhances task completion.
@@
Verification of Relevance Graph Analysis. Our MAID and relevance graph analysis predicts that PSI improves the solvability of MARL learning paradigms (Section 4.2), enabling the vanilla IL learning paradigm more solvable. Experiments verify this across different settings. In simultaneous-move MPE (analysis in Figure 3(c) and results in Figure 5(a)), Independent Q-Learning (IQL) equipped with our PSI achieves task completion comparable to VDN, a CTDE algorithm. This is a notable improvement for IQL over its baseline performance (without our PSI). Similarly, in sequential-move Hanabi (analysis in Figure 4(c) and results in Figure 5(c)), IL algorithms augmented with the PSI achieve performance comparable to, or even better than CTDE algorithms.
Targeted Intervention Outperforming Global Intervention. Our targeted intervention paradigm implemented in PSI consistently outperforms the global intervention paradigm across various MARL backbones, as shown in Figures 5(b) (MPE) and 5(d) (Hanabi). While both paradigms exhibit acyclic relevance graphs suggesting that the corresponding MARL algorithms are solvable in theory, [^5] This can be observed from the MAID analyses in Figures 3(c) and 4(c) for the targeted intervention paradigm, and Figures 3(b) and 4(b) for the global intervention paradigm. the global intervention paradigm’s empirical performance often suffers. We attribute this to the inherent practical difficulty of designing (learning) an effective coordination mechanism to globally assign beneficial, non-conflicting subgoals to multiple agents simultaneously. This can subsequently hinder their learning of optimal independent policies.
@@
Ablation: Our Method vs. Intrinsic Reward. To assess our pre-policy module’s contribution, we conduct an ablation study to compare our PSI against the Intrinsic Reward baseline, both of which belong to the targeted intervention paradigm. Experimental results presented for MPE in Figure 5(a) and for Hanabi in Figure 5(c) reveal a key distinction: while both our method and the Intrinsic Reward baseline effectively lead to the attainment of the additional desired outcome (reflected in the high intrinsic return), our method with the pre-policy module achieves notably superior performance in primary task completion (reflected in the higher extrinsic return). This highlights that the pre-policy module is instrumental for retaining the targeted agent’s adherence to an additional desired outcome while not forgetting the task goal.
@@
Analysis of Nash Equilibrium Convergence in Hanabi. The Hanabi environment, with its multiple distinct equilibria Bard_2020 , serves as an ideal testbed for analyzing convergence to a desired Nash equilibrium. We treat a specific human convention as a desired, high-performing Nash equilibrium (NE) and design the intrinsic reward to directly measure the agents’ compliance to it. As shown in Figure 5(c), the high and stable intrinsic return achieved by our targeted intervention paradigm provides a strong evidence that the agents successfully converge to the desired NE. In contrast, the low intrinsic return of the baseline methods suggest they fail to establish a coordinated convention and are likely stuck in one of Hanabi’s many inferior equilibria.
@@
Appendix I details further exploration of our method’s capabilities and robustness. These include: (1) evaluation using varied additional desired outcomes in MPE and Hanabi; and (2) performance assessments under noisy observation conditions.

### v4

@@ 4 MARL for Sequential Decision Making in MAIDs
We now show how the pre-strategy intervention introduced above is implemented in MARL for sequential decision making. We model MARL using team reward Markov games littman1994markov, and the team reward is represented as the team utility variable. The team utility variable is defined as the total utility variable to identify the preferred Nash equilibrium. The main idea is establishing a connection that a team reward Markov game can be transformed into a multi-agent influence diagram (MAID), allowing our framework to apply. The formal definition of a team reward Markov game and the detailed process of representing it as an MAID by matching their respective variables are provided in Appendix G. Building on this foundation, the subsequent subsections will extend the MARL interaction paradigms introduced in the previous section to team reward Markov games (Section 4.1), discuss the solvability of MARL learning paradigms under these MARL interaction paradigms (Section 4.2), and specify how the pre-strategy intervention in Eq. (3) is integrated into generic MARL algorithms (Section 4.3).
4.1 Extending MARL Interaction Paradigms to Team Reward Markov Games
$D^{t}_{a}$$U_{team}^{t}$$D_{h}^{t}$$S^{t}$$U_{a}^{t}$$U_{h}^{t}$$Z^{t}$$S^{t+1}$
@@ 4.1 Extending MARL Interaction Paradigms to Team Reward Markov Games
Figure 3: MAIDs and their relevance graphs for dynamic simultaneous-move games. For conciseness, we only demonstrate one transition from timesteps $t$ to $t+1$. The variables used in one-shot decision making (see Figure  1) are decorated with timesteps. The semantics of these variables remain unchanged. The pink circles with $s_{t}$ and $s_{t+1}$ indicate states for timesteps $t$ and $t+1$. The green diamonds with $U_{team}^{t}$ indicate the team utility variable that represents the team reward. The incoming edge from an individual utility variable ($U^{t}_{h}$ and $U^{t}_{a}$) to the team utility variable indicates that an agent’s individual contribution to the team reward.
$D^{t}_{a}$$I^{t}$$D^{t}_{h}$$S^{t}$$U^{t}_{a}$$U^{t}_{h}$$Z^{t}$$U_{team}^{t}$$S^{t+1}$
@@
We now demonstrate how to extend the three MARL interaction paradigms (Definition 3.1) to Markov games through the lens of MAIDs. Each interaction paradigm under sequential decision making in dynamic environments with simultaneous or sequential moves can be extended from one-shot decision making introduced in Section 3.2, by considering timesteps and states. Representing the team reward as the team utility variable $U_{team}^{t}$ that is an aggregation of agents’ individual utility variables, we can derive the MAID graphical representation of these three interaction paradigms in team reward Markov games, which are illustrated in Figures 3 and 4.
It can be observed that the widespread global intervention paradigm implemented for MARL algorithms (e.g., applying a centralized coordinator wang2020roma; na2024lagma; liu2023lazy) can be justifiably interpreted with a visualization. More importantly, it opens a new door to design novel MARL interaction paradigms to solve a broad class of problems with the help of MAIDs, such as the targeted intervention paradigm proposed in this paper to solve the team reward Markov game with an additional desired outcome.
@@ 4.2 Solvability of MARL Learning Paradigms under MARL Interaction Paradigms
Independent Learning. Independent learning (IL) (albrecht2024multi)[Ch. 9.3] directly applies single-agent reinforcement learning algorithms to MARL. If the relevance graph is cyclic, individually optimizing each agent’s decision variables using the generalized backward induction algorithm (koller2003multi) cannot guarantee reaching a Nash equilibrium, even when the environmental model (MAIDs) is known—analogous to independent learning in the model-free MARL setting. This aligns with the self-organization paradigm (Figure 3(a) and 4(a)), reflecting the non-stationarity dilemma (hernandez2019survey). The global intervention (Figures 3(b) and 4(b)) and targeted intervention (Figures 3(c) and 4(c)) paradigms can generate acyclic relevance graphs that help mitigate non-stationarity, suggesting that augmenting IL with these two paradigms leads to better performance than its vanilla counterpart, becoming more solvable. [^4] The targeted intervention paradigm can, in principle, be addressed by a mixed learning paradigm with asynchronous updates, guided by the dependency of decision variables in the relevance graph: the targeted agent independently updates its policy, while the other agents update theirs under centralized training in the subsequent turn. Since such an approach is not yet a standard MARL learning paradigm, we leave it for future work.
Centralized Training and Decentralized Execution. As discussed in Remark 2.5, one way to address cyclic relevance graphs is enabling two agents’ decision making matched. One solution is transforming a cyclic relevance graph to a component graph, where each maximal strongly connected component (SCC) is regarded as a supernode de2023graph. More specifically, a maximal SCC consists of all the decision variables that form a cyclic relevance graph at each timestep. Koller and Milch koller2003multi showed that solving the acyclic component graph [^5] A component graph is always acyclic (cormen2022introduction). using the generalized backward induction algorithm can reach a Nash equilibrium when the MAID model is provided. This is associated with the centralized training and decentralized execution (CTDE) (albrecht2024multi)[Ch. 9.1.3] applied in the model-free MARL setting. In summary, CTDE can be seen as an MARL learning paradigm to address the non-stationarity dilemma of vanilla IL under the self-organization paradigm, rather than switching to other MARL interaction paradigms.
@@ 4.3 Pre-Policy Learning in MARL
As we have established the MARL interaction paradigms for team-reward Markov games in MAIDs, the techniques developed in Section 3.3.2 can be applied to realize the pre-strategy intervention by defining the team utility variable $U_{team}^{t}$ as the total utility variable $U_{tot}^{t}$. [^6] By (wang2020shapley), maximizing each agent’s equal contribution to a shared reward is equivalent to maximizing the reward. In the targeted intervention, the team reward is defined as $U_{tot}^{t}=U_{task}^{t}+U_{sec}^{t}$, where $U_{task}^{t}$ is shared among agents and $U_{sec}^{t}$ is only attributed to the targeted agent. Subsequently, the targeted agent’s individual utility variable is defined as $U_{task}^{t}+U_{sec}^{t}$, while other agents’ are defined as $U_{task}^{t}$. Thus, maximizing each agent’s individual utility leads to the preferred NE among multiple equilibria of the primary task utility. Following Definition 3.2, it requires learning a pre-policy $\delta_{pre}:dom(Pa(D))\times\mathcal{Z}\rightarrow\Delta(dom(\sigma_{pre}))$ to generate the pre-strategy $\sigma_{pre}$ for the possible preferred Nash equilibria. This can be realized by maximizing the causal effect expressed in Eq. (3), with replacing the $U_{tot}$ by the cumulative team utilities over $L$ timesteps $\sum_{t=1}^{L}U_{tot}^{t}$, as per Eq. (1). The resulting algorithm, which can be integrated with generic MARL algorithms, is referred to as Pre-Strategy Intervention (PSI).
@@ 5.1 Experimental Setups
We evaluate our proposed Pre-Strategy Intervention (PSI) in two environments: the Multi-Agent Particle Environment (MPE) lowe2017multi, a simultaneous-move cooperative navigation game, and the Hanabi card game Bard_2020, which challenges coordination within a multi-agent system under partial observability. Hanabi is known for having multiple, distinct equilibria, making it an ideal testbed for analyzing convergence to a specific Nash equilibrium. [^7] We choose a convention called “5 Save” for the results in the main paper, while the results of another convention called “The Chop” is reported in Appendix I.2.2. For all scenarios, we select a fixed agent to intervene on, which allows for a controlled analysis of the PSI. Full details of these environments, including specific experimental configurations and how additional desired outcomes are defined and implemented, are provided in Appendix H. In all environments, intrinsic returns are used for measuring reachability of additional desired outcomes, while extrinsic returns are used for measuring primary task completion. Results from 5 random seeds are reported as means with $95\%$ confidence intervals.
Implementation of Our Method. Our PSI is implemented by designing a pre-policy module (a GRU or MLP, matching agent’s backbones), which takes a concatenated input of environmental observations and a measure for an additional desired outcome (formulated as an intrinsic reward). [^8] This implementation maintains solvability of MARL algorithms with the PSI (Appendix F). It outputs an embedding, which is then fed to Q-value functions for value-based MARL methods or critics for policy-based methods.
Baselines. All baselines are implemented with the same architecture and training setups as our method. We evaluate our method against several base MARL algorithms referred to as Base MARL (self-organization paradigms): for MPE, these include IQL (tan1993multi), VDN (sunehag2017value) and QMIX (rashid2020monotonic); for Hanabi, we use IPPO (schulman2017proximal), MAPPO (yu2022surprising) and PQN (gallici2024simplifying) (with PQN-IQL for IL and PQN-VDN for CTDE variants). An ablation study that removes the pre-policy module while retaining the intrinsic reward maximization, referred to as Intrinsic Reward, examines the effectiveness of the pre-policy module. To emphasize the strength of our targeted intervention paradigm, we also compare PSI against the global intervention version of PSI, referred to as GPSI, which applies the pre-policy module to intervene on all agents. Finally, we include baselines of the global intervention paradigm that only focuses on primary task completion, such as the modified versions of LIIR du2019liir and LAIES liu2023lazy, as representative policy-based and value-based methods, respectively.
@@ 5.2 Results and Analysis
(a) MPE: Main results.
(b) Hanabi: Main results.
(c) Hanabi: PSI vs. GPSI.
(d) PSI vs. Additional-Outcome-Free Approaches.
@@
Figure 5: In (a) and (b), PSI is compared against Intrinsic Reward and Base MARL; these plots show both extrinsic (the primary task completion) and intrinsic (reachability of an additional desired outcome) returns. In (c), the targeted intervention paradigm (PSI) is compared against the global intervention paradigm (GPSI). In (d), PSI is compared against approaches that only focus on the primary task completion with no consideration of any additional desired outcome.
Coordination Achieving When Intervening on a Single, Targeted Agent. As Figure 5(d) shows, our PSI can outperform both LIIR and LAIES on the primary task completion in both MPE and Hanabi. This result confirms that coordination can be achieved when assigning an additional desired outcome (e.g. convention) derived from human knowledge, to a single targeted agent through the targeted intervention paradigm, answering our motivating research question. Moreover, this may even improve primary task completion, compared with approaches of the global intervention paradigm that does not consider any additional desired outcome, avoiding the issue of miscoordination.
@@
Verification of Relevance Graph Analysis. Our MAID and relevance graph analysis predicts that PSI improves the solvability of MARL learning paradigms (Section 4.2), enabling the vanilla IL learning paradigm more solvable. Experiments verify this across different settings. In simultaneous-move MPE (analysis in Figure 3(c) and results in Figure 5(a)), IQL equipped with our PSI achieves task completion comparable to VDN, a CTDE algorithm. This is a notable improvement for IQL over its baseline performance (without our PSI). Similarly, in sequential-move Hanabi (analysis in Figure 4(c) and results in Figure 5(b)), IL algorithms augmented with the PSI also achieve performance comparable to, or even better than CTDE algorithms.
Targeted Intervention vs. Global Intervention. Our targeted intervention paradigm (PSI) consistently outperforms the global intervention paradigm (GPSI), as shown in Figures 5(c). Although both paradigms achieve comparable performance on the primary task completion, the global intervention paradigm often struggles to reach the additional desired outcome. We attribute this to the inherent practical difficulty of designing (learning) an effective coordination mechanism to globally assign beneficial, non-conflicting goals to multiple agents simultaneously. This highlights the potential application of PSI to safety-critical systems (wang2021multi), which require strictly controlled operation.
@@
Ablation Study on the Pre-Policy Module. To assess our pre-policy module’s contribution, we conduct an ablation study to compare PSI against the Intrinsic Reward baseline. Experimental results presented for MPE in Figure 5(a) and for Hanabi in Figure 5(b) reveal a key distinction: while both PSI and the Intrinsic Reward baseline effectively lead to the attainment of the additional desired outcome, the PSI (with the pre-policy module) achieves notably superior performance on the primary task completion. This highlights that the pre-policy module is instrumental for retaining the targeted agent’s adherence to an additional desired outcome while not forgetting the primary task goal.
@@
Analysis of Nash Equilibrium Convergence in Hanabi. The Hanabi environment, with its multiple distinct equilibria Bard_2020, serves as an ideal testbed for analyzing convergence to a preferred Nash equilibrium (NE). We treat a specific human convention as a preferred, high-performing NE and design the intrinsic reward to directly measure the agents’ compliance to it. As shown in Figure 5(b), the high and stable intrinsic return achieved by our PSI provides a strong evidence that the agents successfully converge to the preferred NE. In contrast, the low intrinsic return of the baselines suggest they fail to establish the convention and are likely stuck in one of Hanabi’s many inferior equilibria.
@@
Additional Results. Appendix I details further exploration of our method’s capabilities and robustness. These include: (1) evaluation using varied additional desired outcomes in MPE and Hanabi; and (2) performance assessments under noisy observation conditions.